Tantalum is a rare, hard, blue-gray, lustrous transition metal that is highly corrosion-resistant. It has a density of 16.4 grams per cubic centimeter and an atomic weight of 180.948 g/mol. What is the atomic radius of Ta if it adopts the body centered cubic unit structure? Express your answer in Å.


1
Expert's answer
2021-09-01T03:11:55-0400

16/180.948 = 0.0884

6.023 × 1023 ×0.0884 = 5.33 × 1022 A